# Local Dissipativity Analysis of Nonlinear Systems

### v1

@@ 3.2 Quadratic Error Bounds
Section 3.1 bounded matrix behavior across a simplex of a triangulation. Here, the focus shifts to bounding an inequality constraint within a closed two-norm ball about the origin. As in previous sections, application of Taylor’s theorem is used to bound the inequality across the region. However, by focusing on a two-norm ball about the origin, Theorem 11 later expands concepts used in CPA function synthesis to different function types.
@@ 4 Local dissipativity analysis
The error bounds developed in Section 3 are now leveraged to formulate conditions for which a CPA function or combined CPA and quadratic function is the storage function of a dissipative system. These results are then leveraged in Section 4.2 to develop convex optimization problems that simultaneously synthesize a system’s storage function and analyzes common QSR characterizations – such as gain, passivity, and conic bounds.
@@ 4.1 Storage function conditions
Theorems 12 and 13 give conditions for a CPA function or a combined CPA and quadratic function, respectively, adhere to the QSR dissipativity condition (5) across a region of the state space. Theorem 13 considers the case where the system dynamics are such that any input affine term disappears at the origin, while Theorem 13 allows for non-zero control affine terms at the origin. This distinction is important, as all error bounds in the optimization problem must be zero at the origin to prevent infeasibility. While this always occurs when the input term disappears, a modified quadratic and CPA storage function (which is quadratic at the origin) is needed for the nonzero case.
@@
Consider (1) with $\mathbf{B}=0$ and $\mathbf{D}=0$. Suppose that $f,G,h,J\in\mathcal{C}^{2}$ on the set $\mathcal{X}\in\mathfrak{R}^{n}$. Let $\mathbf{Q}=\mathbf{Q}^{\top}\in\mathbb{R}^{p\times p}$, $\mathbf{S}\in\mathbb{R}^{p\times m}$, and $\mathbf{R}=\mathbf{R}^{\top}\in\mathbb{R}^{m\times m},$ where $\mathbf{Q}\prec 0$ and $\mathbf{R}\succ 0$. Let $\Pi_{1}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3}\in\mathbb{R}^{p\times p}$ be diagonal, positive definite matrices, where $\underline{\pi}_{z}$ is the minimum diagonal value of $\Pi_{z}^{-1}$. Define $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}$. Define the CPA function, $V=\{V_{x}\}_{x\in\mathbb{E}_{\mathcal{T}}}$. Consider
$V_{x}$ $\geq 0\quad\forall x\in\mathbb{E}_{x}$ (14a)
@@
where $\hat{\mathbf{R}}=\frac{1}{2}(\Pi_{1}^{-1}{+}\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I})$, $\hat{\mathbf{Q}}=\frac{1}{2}(\Pi_{2}^{-1}{+}\Pi_{3}^{-1})$, $\hat{f}_{i,j}=\frac{n}{2}l\beta_{i}c_{i,j}$, $\hat{G}_{i,j}=n\begin{bmatrix}\mu_{1}c_{i,j}&\ldots&\mu_{m}c_{i,j}\end{bmatrix}^{\top}$, $\hat{h}_{i,j}=n\begin{bmatrix}\rho_{1}c_{i,j}&\ldots&\rho_{p}c_{i,j}\end{bmatrix}^{\top}$, and $\hat{J}_{i,j}=n\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I}.$ Further, $c_{i,j}$ is defined by (9) when $0\in\sigma$ and by (33) when $0\notin\sigma,$
$\beta_{i}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial f^{(p)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$, (16)
$\rho_{i,a}\geq\max_{q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial h^{(a)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$, (17)
$\mu_{i,k}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial G_{k}^{(p)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$, (18)
@@
$\theta_{i}\geq\max_{k\in\mathbb{Z}_{1}^{m},a\in\mathbb{Z}_{1}^{p},q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial J_{k}^{(a)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$. (19)
If $V$ satisfies (14), then $V$ satisfies the dissipation HJI (5) for all $x\in\mathcal{X}^{\circ}$ for $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
@@
$\frac{1}{2}\left\lVert\nabla V\right\rVert_{1}\sum_{j=0}^{n}\!\lambda_{j}\left\lVert\overline{\mathbf{H}}_{f}(x_{i,j},x)\right\rVert_{\infty}$. From here, Lemma 17 is be applied. To bound $\frac{1}{2}\sum_{j=0}^{n}\!\lambda_{j}(\frac{1}{2}\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)-\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\mathbf{S})\Pi_{1}(\frac{1}{2}\overline{\mathbf{H}}_{G}(x_{i,j},x)^{\top}\nabla V-\mathbf{S}\overline{\mathbf{H}}_{h}(x_{i,j},x)),$ apply the Cauchy-Schwarz inequality, the triangle inequality, and the properties of matrix operator norms to get $\frac{1}{2}\left\lVert\Pi\right\rVert_{1}\sum_{j=0}^{n}\!\lambda_{j}\Bigl(\left\lVert\overline{\mathbf{H}}_{h}(x_{i,j},x)\right\rVert_{2}^{2}\left\lVert\mathbf{S}\right\rVert_{2}^{2}+\frac{1}{2}\left\lVert\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)\right\rVert_{2}^{2}\Bigr).$ Note also that
$\left\lVert\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)\right\rVert_{2}^{2}\leq\left\lVert\nabla V\right\rVert_{1}^{2}\sum_{k=1}^{m}\left\lVert\overline{\mathbf{H}}_{G_{k}}(x_{i,j},x)\right\rVert_{\infty}^{2}$. From here, Lemma 9 is applied to terms involving $h$ and $G.$ Finally, $\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\Pi_{2}\overline{\mathbf{H}}_{h}(x_{i,j},x)=\sum_{k=1}^{p}\overline{\mathbf{H}}_{h^{(k)}}(x_{i,j},x)^{2}\pi_{2,k}$ can be bounded using the same procedure as in Lemma 17. Therefore,
@@
By assumption, (14) holds on each $n$-simplex of the $\mathcal{T}$. By Theorem 10, (5) holds for all $\mathbf{x}\in\mathcal{X}^{\circ}$ for some $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
@@
Consider (1). Suppose that $f,G,h,J\in\mathcal{C}^{2}$ on the set $\mathcal{X}\in\mathfrak{R}^{n}.$ Let $\mathbf{Q}=\mathbf{Q}^{\top}\in\mathbb{R}^{p\times p}$, $\mathbf{S}\in\mathbb{R}^{p\times m}$, and $\mathbf{R}=\mathbf{R}^{\top}\in\mathbb{R}^{m\times m},$ where $\mathbf{Q}\prec 0$ and $\mathbf{R}\succ 0$. Let $\Pi_{1},\Pi_{4},\Pi_{6}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3},\Pi_{5},\Pi_{7},\Pi_{8},\Pi_{9}\in\mathbb{R}^{p\times p}$ be diagonal, positive definite matrices, where $\underline{\pi}_{z}$ is the minimum diagonal value of a matrix $\Pi_{z}^{-1}$. Consider the sets $0\in\Psi$ and $B_{\epsilon}(0)$, where $\Psi\subset B_{\epsilon}(0)\subset\mathcal{X}.$ Define a triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}$ on $\mathcal{X}\setminus\Psi^{\circ}$. Define a CPA function $V_{C}=\{\bar{V}_{x}\}_{x\in\mathbb{E}_{\mathcal{T}}}$ and a quadratic function $V_{P}=x^{\top}\mathbf{P}x$. Let there exist a function
[…]
Consider
@@
$\left\lVert\mathbf{S}\right\rVert_{2}$ $\leq\hat{s}$, (21e)
$\mathbf{M}_{\epsilon}$ $\preceq 0,$ (21f)
$\bar{V}_{x}$ $\geq 0\quad\forall x\in\mathbb{E}_{\mathcal{T}}$ (21g)
$\left\lVert\nabla\bar{V}_{i}\right\rVert_{1}$ $\leq l_{i}\quad\forall i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},$ (21h)
$\mathbf{M}(x_{i,j})$ $\preceq 0,\quad\forall i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},\forall j\in\mathbb{Z}_{0}^{n},$ (21i)
where $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}\in\mathbb{R},$ $\mathbf{M}$ is defined by (15) and
$\mathbf{M}_{\epsilon}=$ (22)
@@
$\beta_{\epsilon}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in B_{\epsilon}(0)}\left\lvert\frac{\partial^{2}f^{(p)}(\mathbf{x})}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Bigr|_{\mathbf{x}=\xi}\right\rvert$, (24)
@@
Further, $\beta_{i},\rho_{i,a}$ and $\mu_{i,k}$ are defined by (16), (17), and (18), and $c_{i,j}$ is defined by (33).
If $V$ satisfies (21), then $V$ satisfies the dissipation HJI (2) for all $x\in\mathcal{X}^{\circ}$ for $\mathbf{Q}$, $\mathbf{S}$, and $\mathbf{R}.$
@@
This proof first shows that (2) holds for $V_{C}$ in $\mathcal{X}^{\circ}\setminus\Psi$ and $V_{P}$ in $B_{\epsilon}(0).$ Then, it is shown to hold for $V$ in $\mathcal{X}^{\circ}.$
When $V_{C}, VV$ is a CPA function and the proof follows similarly to Theorem 12 to show that the dissipation HJI holds for $\mathbf{Q},$ $\mathbf{S},$ and $\mathbf{R}$ on $\mathcal{X}\setminus\Psi^{\circ}$ with function $V_{C}$. When $V=V_{P},$ V is quadratic, Theorem 11 must be leveraged to ensure that the dissipation HJI holds on $B_{\epsilon}(0)$.
Constraint (22) creates an upper bound on the QSR dissipativity LMI (5) for all $\mathbf{x}\in B_{\epsilon}(0)$ using Theorem 11. To develop this constraint, perform a Taylor series expansion on each function within the QSR inequality about the origin using $\nabla V=\mathbf{x}^{\top}\mathbf{P}+\mathbf{P}\mathbf{x}$, i.e. $\mathbf{x}^{\top}\mathbf{P}\hat{f}+\hat{f}^{\top}\mathbf{P}\mathbf{x}-\hat{h}^{\top}\mathbf{Q}\hat{h}+(\mathbf{x}^{\top}\mathbf{P}\hat{G}-\hat{h}^{\top}\mathbf{Q}\hat{J}-\hat{h}^{\top}\mathbf{S})\hat{\mathbf{R}}^{-1}(\hat{G}^{\top}\mathbf{P}\mathbf{x}-\hat{J}^{\top}\mathbf{Q}\hat{h}-\mathbf{S}^{\top}\hat{h})\leq 0,$ where $\hat{\mathbf{R}}=\mathbf{R}+\hat{J}^{\top}\mathbf{S}+\mathbf{S}^{\top}\hat{J}+\hat{h}^{\top}\mathbf{Q}\hat{h}$. For example,
$\hat{G}$= $\mathbf{B}+\tilde{G}(\mathbf{x})=\mathbf{B}+\begin{bmatrix}\mathbf{J}_{g_{1}}(0)\mathbf{x}&\ldots\mathbf{J}_{g_{m}}(0)\mathbf{x}\end{bmatrix}$+
$\frac{1}{2}\begin{bmatrix}\begin{bmatrix}\mathbf{x}^{\top}\mathbf{H}_{g^{(1)}_{1}}(\mathbf{z}_{g^{(1)}_{1}})\\ \vdots\\ \mathbf{x}^{\top}\mathbf{H}_{g^{(n)}_{1}}(\mathbf{z}_{g^{(n)}_{1}})\end{bmatrix}\mathbf{x}&\ldots&\begin{bmatrix}\mathbf{x}^{\top}\mathbf{H}_{g^{(1)}_{m}}(\mathbf{z}_{g^{(1)}_{m}})\\ \vdots\\ \mathbf{x}^{\top}\mathbf{H}_{g^{(n)}_{m}}(\mathbf{z}_{g^{(n)}_{m}})\end{bmatrix}\mathbf{x}\end{bmatrix}$,
where the Hessian of each element is evaluated at a convex combination of $\mathbf{x}$ and 0.
Factor $\mathbf{x}$ out on either side of the inequality. Then a Schur complement is performed about terms involving $\hat{\mathbf{R}}$ and another Schur complement is performed about terms involving $\mathbf{Q}.$ Apply the process of Theorem 11. Note that any term on the diagonal of the matrix that contains $\mathbf{x}$ is bounded above using the matrix two norm, e.g. $\mathbf{S}^{\top}\hat{J}\leq\left\lVert\mathbf{S}\right\rVert_{2}\left\lVert\hat{J}\right\rVert_{2}.$ The resulting LMI is
@@
From Theorem 11, (5) holds for all $\mathbf{x}\in B_{\epsilon}(0)$ with $V_{P}$ for $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
@@
Consider (1). Let $\Omega\subseteq\mathcal{X}\in\mathfrak{R}^{n}$ be control invariant for the signals $\mathcal{U}_{\Omega}\subseteq\mathcal{L}_{2}^{m}$. If Theorem 12 or Theorem 13 hold over $\mathcal{X},$ then the family of mappings $\mathcal{G}_{\Omega}$ is locally QSR dissipative in $\Omega.$
@@
By assumption, the QSR HJI (5) holds for all $x\in\mathcal{X}$ and all $\mathbf{u}\in\mathcal{L}_{2}^{m}.$ Thus, (5) also holds for all $x\in\Omega$ and $\mathbf{u}\in\mathcal{U}_{\Omega}$ and the mappings $\mathcal{G}_{\Omega}(x_{0})$ are locally dissipative for all $x_{0}\in\Omega$.

### v2

@@ 3.2 Quadratic Error Bounds
Section 3.1 bounded matrix behavior across a simplex of a triangulation. Here, the focus shifts to bounding an inequality constraint within a closed two-norm ball about the origin. As in previous sections, Taylor’s Theorem bounds the inequality across the region. However, by focusing on a two-norm ball about the origin, Theorem 11 expands concepts used in CPA function synthesis to different function types.
@@ 4 Local dissipativity analysis
The error bounds developed in Section 3 are now leveraged to formulate conditions for which a CPA function or combined CPA and quadratic function is the storage function of a dissipative system. Section 4.2 then uses these to develop convex optimization problems that simultaneously synthesize a system’s storage function and analyze common QSR characterizations – such as gain, passivity, and conic bounds.
@@ 4.1 Storage function conditions
Theorems 12 and 13 give conditions for a CPA function or a combined CPA and quadratic function, respectively, adhere to the QSR dissipativity condition (5) across a region of the state space. Theorem 12 considers the case where the system dynamics make all input affine terms disappear at the origin, while Theorem 13 admits non-zero control affine terms at the origin. This distinction is important, as all error bounds in the optimization problem must be zero at the origin to prevent infeasibility. While this always occurs when the input term disappears, a modified quadratic and CPA storage function (which is quadratic at the origin) is needed for the nonzero case.
@@
Consider (1) with $\mathbf{B}=0$ and $\mathbf{D}=0$. Suppose that $f,G,h,J\in\mathcal{C}^{2}$ on the set $\mathcal{X}\in\mathfrak{R}^{n}$ triangulated by $\mathcal{T}.$ Let $\mathbf{Q}=\mathbf{Q}^{\top}\in\mathbb{R}^{p\times p}$, $\mathbf{S}\in\mathbb{R}^{p\times m}$, and $\mathbf{R}=\mathbf{R}^{\top}\in\mathbb{R}^{m\times m},$ where $\mathbf{Q}\prec 0$ and $\mathbf{R}\succ 0$. Let $\Pi_{1}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3}\in\mathbb{R}^{p\times p}$ be diagonal, positive definite matrices, where $\underline{\pi}_{z}$ is the minimum diagonal value of $\Pi_{z}^{-1}$. Define $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}$. Define the CPA function, $V=\{V_{x}\}_{x\in\mathbb{E}_{\mathcal{T}}}$. Consider
$V_{x}$ $\geq 0\quad\forall x\in\mathbb{E}_{\mathcal{T}}$ (14a)
@@
where $\hat{\mathbf{R}}=\frac{1}{2}(\Pi_{1}^{-1}{+}\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I})$, $\hat{\mathbf{Q}}=\frac{1}{2}(\Pi_{2}^{-1}{+}\Pi_{3}^{-1})$, $\hat{f}_{i,j}=\frac{n}{2}l\beta_{i}c_{i,j}$, $\hat{G}_{i,j}=n\begin{bmatrix}\mu_{1}c_{i,j}&\ldots&\mu_{m}c_{i,j}\end{bmatrix}^{\top}$, $\hat{h}_{i,j}=n\begin{bmatrix}\rho_{1}c_{i,j}&\ldots&\rho_{p}c_{i,j}\end{bmatrix}^{\top}$, and $\hat{J}_{i,j}=n\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I}.$ Further, $c_{i,j}$ equals $\overline{c}_{j}$ in (9) when $0\in\sigma$ and follows (32) when $0\notin\sigma,$
$\beta_{i}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial f^{(p)}}{\partial x^{(q)}\partial x^{(r)}}\Big|_{x=\xi}\right\rvert$, (16)
$\rho_{i,a}\geq\max_{q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial h^{(a)}}{\partial x^{(q)}\partial x^{(r)}}\Big|_{x=\xi}\right\rvert$, (17)
$\mu_{i,k}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial G_{k}^{(p)}}{\partial x^{(q)}\partial x^{(r)}}\Big|_{x=\xi}\right\rvert$, (18)
@@
$\theta_{i}\geq\max_{k\in\mathbb{Z}_{1}^{m},a\in\mathbb{Z}_{1}^{p},q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial J_{k}^{(a)}}{\partial x^{(q)}\partial x^{(r)}}\Big|_{x=\xi}\right\rvert$. (19)
If $V$ satisfies (14), then $V$ satisfies the dissipation condition (5) for all $x\in\mathcal{X}^{\circ}$ for $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
@@
$\frac{1}{2}\left\lVert\nabla V\right\rVert_{1}\sum_{j=0}^{n}\!\lambda_{j}\left\lVert\overline{\mathbf{H}}_{f}(x_{i,j},x)\right\rVert_{\infty}$. From here, Lemma 18 is be applied. To bound $\frac{1}{2}\sum_{j=0}^{n}\!\lambda_{j}(\frac{1}{2}\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},\!x)$
$\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\mathbf{S})\Pi_{1}(\frac{1}{2}\overline{\mathbf{H}}_{G}(x_{i,j},x)^{\top}\nabla V-\mathbf{S}\overline{\mathbf{H}}_{h}(x_{i,j},x)),$ apply the Cauchy-Schwarz inequality, the triangle inequality, and the properties of matrix operator norms to get $\frac{1}{2}\left\lVert\Pi\right\rVert_{1}\sum_{j=0}^{n}\!\lambda_{j}(\left\lVert\overline{\mathbf{H}}_{h}(x_{i,j},x)\right\rVert_{2}^{2}\left\lVert\mathbf{S}\right\rVert_{2}^{2}$
$\frac{1}{2}\left\lVert\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)\right\rVert_{2}^{2}).$ Note also that
$\left\lVert\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)\right\rVert_{2}^{2}\leq\left\lVert\nabla V\right\rVert_{1}^{2}\sum_{k=1}^{m}\left\lVert\overline{\mathbf{H}}_{G_{k}}(x_{i,j},x)\right\rVert_{\infty}^{2}$. From here, Lemma 9 is applied to terms involving $h$ and $G.$ Finally, $\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\Pi_{2}\overline{\mathbf{H}}_{h}(x_{i,j},x)$
$\sum_{k=1}^{p}\overline{\mathbf{H}}_{h^{(k)}}(x_{i,j},x)^{2}\pi_{2,k}$ can be bounded using the same procedure as in Lemma 18. Therefore,
@@
By assumption, (14) holds on each $n$-simplex of the $\mathcal{T}$. By Theorem 10, (5) holds for all $x\in\mathcal{X}^{\circ}$ for some $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
@@
Consider (1). Suppose that $f,G,h,J\in\mathcal{C}^{2}$ on the set $\mathcal{X}\in\mathfrak{R}^{n}.$ Let $\mathbf{Q}=\mathbf{Q}^{\top}\in\mathbb{R}^{p\times p}$, $\mathbf{S}\in\mathbb{R}^{p\times m}$, and $\mathbf{R}=\mathbf{R}^{\top}\in\mathbb{R}^{m\times m},$ where $\mathbf{Q}\prec 0$ and $\mathbf{R}\succ 0$. Let $\Pi_{1},\Pi_{4},\Pi_{5},\Pi_{6}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3},\Pi_{7},\Pi_{8},\Pi_{9}\in\mathbb{R}^{p\times p}$ be diagonal, positive definite matrices, where $\underline{\pi}_{z}$ is the minimum diagonal value of a matrix $\Pi_{z}^{-1}$. Consider the sets $0\in\Psi$ and $B_{\epsilon}(0)$, where $\Psi\subset B_{\epsilon}(0)\subset\mathcal{X}.$ Define a triangulation $\mathcal{T}{=}\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}$ on $\mathcal{X}\setminus\Psi^{\circ}$. Let there exist a CPA function $V_{C}{=}\{\bar{V}_{x}\}_{x\in\mathbb{E}_{\mathcal{T}}}$ satisfying (14) and a quadratic function $V_{P}=x^{\top}\mathbf{P}x$ satisfying
@@
$\mathbf{M}_{\epsilon}$ $\preceq 0,$ (20e)
where
$\mathbf{M}_{\epsilon}=$
@@
$\beta_{\epsilon}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in B_{\epsilon}(0)}\left\lvert\frac{\partial^{2}f^{(p)}(x)}{\partial x^{(q)}\partial x^{(r)}}\Bigr|_{x=\xi}\right\rvert$, (22)
@@
Further, $\beta_{i},\rho_{i,a}$ and $\mu_{i,k}$ are defined by (16), (17), and (18), and $c_{i,j}$ is defined by (32). Then
[…]
satisfies the dissipation condition (2) for all $x\in\mathcal{X}^{\circ}$ for $\mathbf{Q}$, $\mathbf{S}$, and $\mathbf{R}.$
@@
On the interior of $\{x\in\mathcal{X}\mid V(x)=V_{C}(x)\}$, $V=V_{C}$ is a CPA function and the proof follows similarly to Theorem 12. On the interior of $\{x\in\mathcal{X}\mid V(x)=V_{P}(x)\}$, $V=V_{P}$ is quadratic, so Theorem 11 must be leveraged to ensure that the dissipation condition holds on $B_{\epsilon}(0)$. Lastly, we will establish continuity of $V$ so that its Dini Derivative is well-defined on $\mathcal{X}$, meaning (2) inherits satisfaction on the boundary boundary where $V(x)=V_{C}(x)=V_{P}(x)$.
Constraint (13) creates an upper bound on the QSR dissipativity condition (5) for all $x\in B_{\epsilon}(0)$ using Theorem 11. To develop this constraint, perform a Taylor series expansion on each function within the QSR inequality about the origin using $\nabla V=x^{\top}\mathbf{P}+\mathbf{P}x$, i.e. $x^{\top}\mathbf{P}\hat{f}+\hat{f}^{\top}\mathbf{P}x-\hat{h}^{\top}\mathbf{Q}\hat{h}+(x^{\top}\mathbf{P}\hat{G}-\hat{h}^{\top}\mathbf{Q}\hat{J}-\hat{h}^{\top}\mathbf{S})\hat{\mathbf{R}}^{-1}(\hat{G}^{\top}\mathbf{P}x-\hat{J}^{\top}\mathbf{Q}\hat{h}-\mathbf{S}^{\top}\hat{h})\leq 0,$ where $\hat{\mathbf{R}}=\mathbf{R}+\hat{J}^{\top}\mathbf{S}+\mathbf{S}^{\top}\hat{J}+\hat{h}^{\top}\mathbf{Q}\hat{h}$. For example,
$\hat{G}$= $\mathbf{B}+\tilde{G}(x)=\mathbf{B}+\begin{bmatrix}\mathbf{J}_{g_{1}}(0)x&\ldots\mathbf{J}_{g_{m}}(0)x\end{bmatrix}$+
$\frac{1}{2}\begin{bmatrix}\begin{bmatrix}x^{\top}\mathbf{H}_{g^{(1)}_{1}}(\mathbf{z}_{g^{(1)}_{1}})\\ \vdots\\ x^{\top}\mathbf{H}_{g^{(n)}_{1}}(\mathbf{z}_{g^{(n)}_{1}})\end{bmatrix}x&\ldots&\begin{bmatrix}x^{\top}\mathbf{H}_{g^{(1)}_{m}}(\mathbf{z}_{g^{(1)}_{m}})\\ \vdots\\ x^{\top}\mathbf{H}_{g^{(n)}_{m}}(\mathbf{z}_{g^{(n)}_{m}})\end{bmatrix}x\end{bmatrix}$,
where the Hessian of each element is evaluated at a convex combination of $x$ and 0.
Factor $\mathbf{x}$ out on either side of the inequality. Then a Schur complement is performed about terms involving $\hat{\mathbf{R}}$ and another Schur complement is performed about terms involving $\mathbf{Q}.$ Apply the process of Theorem 11. Note that any term on the diagonal of the matrix that contains $x$ is bounded above using the matrix two norm, e.g. $\mathbf{S}^{\top}\hat{J}\leq\left\lVert\mathbf{S}\right\rVert_{2}\left\lVert\hat{J}\right\rVert_{2}.$ The resulting LMI is
@@
From Theorem 11, (5) holds for all $x\in B_{\epsilon}(0)$ with $V_{P}$ for $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
@@
For dynamics (1) and set $\Omega\subseteq\mathcal{X}\in\mathfrak{R}^{n}$, let $\mathcal{U}_{\Omega}\subseteq\mathcal{L}_{2}^{m}$ contain a neighborhood of the origin. If Theorem 12 or Theorem 13 holds over $\mathcal{X},$ then the family of mappings $\mathcal{G}_{\Omega}$ is locally QSR dissipative in $\Omega.$
@@
By assumption, the QSR condition (5) holds for all $x\in\mathcal{X}$. Thus, (5) also holds for all $x\in\Omega$ and $\mathbf{u}\in\mathcal{U}_{\Omega}$ and the mappings $\mathcal{G}_{\Omega}(x_{0})$ are locally dissipative for all $x_{0}\in\Omega$.